OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step. 1 - (1*2) + -1 2

OpenStudy (anonymous):

Simplify: \[ 1-(1*2)+-1 \]?

OpenStudy (anonymous):

no its one half times 1 times two plus negative 1

OpenStudy (anonymous):

All right: \[ \frac{1}{2}(1\cdot 2)+-1 \] Multiplicative identity: \[ \frac{1}{2}(1\cdot 2)+-1=\frac{1}{2}(2)+-1 \]Definition of Multiplicative Inverse: \[ \frac{1}{2}(2)+-1=1+-1 \]Definition of Additive inverse: \[ 1+-1=0 \]

OpenStudy (anonymous):

thank you!!!!! ur seriously a genius